# A transversality theorem for semi-algebraic sets with application to signal recovery from the second moment and cryo-EM

### v2

@@ 3 Signal recovery from the second moment
Most applications we consider in this paper involve signal recovery from the second moment. Thus, we are interested in the transversality of semi-algebraic sets with orbits of the ambiguity group of the second moment. In this section, we first present the model of recovering a signal translated by random group actions from the second moment and then specialize Theorem 2.2 and Theorem 2.3 to this case. Finally discuss two important applications: factoring Gram matrices and phase retrieval.
@@ 3.2 Main results for signal recovery from the second moment
We view $V$ as a representation of $H$, where the $\ell$-th component of $H$, $\mathrm{O}(N_{\ell})$, acts diagonally. This means that if we view an element of $V_{\ell}^{R_{\ell}}$ as a tuple $v_{\ell}=(v_{\ell}[1],v_{\ell}[2],\ldots,v_{\ell}[R_{\ell}])$ with $v_{\ell}[r]\in V_{\ell}$, then $g\cdot v_{\ell}=(g\cdot v_{\ell}[1],g\cdot v_{\ell}[2],\ldots g\cdot v_{\ell}[R_{\ell}])$. Then, the orbits of $H$ have dimension
@@
In Section 4 we discuss the implications of Corollary 3.1 to cryo-EM—the prime motivation of this paper. Before we do this we briefly discuss two additional applications: factoring Gram matrices under semi-algebraic priors and phase retrieval.
@@ 3.3.1 Factoring Gram matrices under semi-algebraic priors
As a warmup application, we consider the simple case above where $L=1$; i.e., $V=W^{\oplus R}$ with $W$ an $N$-dimensional irreducible representation of $H$. An element $X\in V$ is simply an $N\times R$ matrix, and the second moment is the Gram matrix $X^{T}X$. Generally, the Gram matrix determines $X$ up to multiplication by a matrix in $O(N)$. However, Corollary 3.1 implies that an $N\times R$ matrix can be determined from its Gram matrix if it lies in a suitably low-dimensional generic semi-algebraic subset of $\mathbb{R}^{N\times R}$.
@@
    If $X$ is a generic matrix in an $\mathrm{O}(V)$-generic semi-algebraic subset of dimension $M<NR-\left(\dim\mathrm{O}(N)-\dim\mathrm{O}(N-R)\right)-2$, then $X$ can be recovered (up to a sign for $\mathrm{GL}(V)$) from its Gram matrix. Likewise, if $22M<NR-\left(\dim\mathrm{O}(N)-\dim\mathrm{O}(N-R)\right)-$ 2, then every $X$ in an $\mathrm{O}(V)$-generic semi-algebraic subset can be recovered up to a sign from its Gram matrix.
@@ 3.3.2 Phase retrieval
The phase retrieval problem entails recovering a signal from its power spectrum: the magnitudes of its Fourier transform. Clearly, the power spectrum determines the signal up to a product of $N$ phases (unitary matrices of dimension 1) in Fourier space. This problem has numerous applications in optical imaging, X-ray crystallography and more, see [43, 10, 27, 31], and references therein. Semi-algebraic priors are essential in phase retrieval. For example, in X-ray crystallography, the goal is to recover the sparsely-spread atoms that from an image [26, 13, 14]. In addition, deep generative models have been proven effective in different branches of phase retrieval in recent years [44, 40, 35, 32, 21].
Explicitly, the goal in phase retrieval is recovering a signal $x\in\mathbb{R}^{N}$ from its power spectrum,
@@
where $\hat{x}[i]$ is the $i$-th entry of discrete Fourier transform of the signal $x$. The analysis of [24, Section 6.2] shows that the power spectrum is equivalent to the second moment of $x$ with respect to the action of the dihedral group $D_{2N}$, which acts on $\mathbb{R}^{N}$ by translation and reflection. As a representation of $D_{2N}$, the vector space $\mathbb{R}^{N}$ decomposes as a sum of irreducibles as follows.
@@
    $\mathbb{R}^{N}=V_{0}\oplus V_{1}\oplus\ldots V_{N/2-1}\oplus V_{N/2},$
@@
    $\mathbb{R}^{N}=V_{0}\oplus V_{1}\oplus\ldots V_{(N-1)/2},$
@@
In particular, the dimension of the orbits of $H=\prod\mathrm{O}(V_{i})$ is $N/2-1$ if $N$ is even and $(N-1)/2$ if $N$ is odd. Thus, we can invoke Corollary 3.1 with $K=N/2+1$ ($N$ even) and $K=(N+1)/2$ ($N$ odd) to obtain the following result.
@@
In [24], the authors considered phase retrieval for sparse vectors in generic bases. There, using algebro-geometric arguments, the bounds $N\geq 2M-1$ (generic signal recovery) and $N\geq 4M-3$ (all signal recovery) are obtained.
@@ 4 Cryo-EM
[…]
One of the main challenges in processing cryo-EM data sets, and analyzing their performance, is the extremely high noise level. This is due to the limited number of electrons the microscope can transmit without damaging the biological sample. From an information-theoretic perspective, it was shown that in the high-noise regime $\sigma\to\infty$, the minimal number of samples required for accurate recovery, regardless of any specific algorithm, needs to scale as $n/\sigma^{2d}\to\infty$, where $d$ is the lowest order moment of the observations that determine the sought structure uniquely [2] [^3] This result is true when the dimension of the signal is finite, as we assume in this paper. In the very high-dimensional regime, the sample complexity is governed by other factors [41].. In [7], the authors gave computational evidence that the third-order moment determines generic signals while the second moment does not, implying that the minimal number of observations $n$ for generic signals scales rapidly with the noise level as $\sigma^{6}$. This motivates identifying classes of signals that can be determined from the second moment, and thus with fewer samples. For example, recent works showed that if the structure can be represented with only a few coefficients under some basis [15], or as a sparse mixture of Gaussians [16], then the signal can be recovered from the second moment. In this work, we consider the more general family of signals that lie in a semi-algebraic set (including sparse signals as in [15]) and derive conditions under which a molecular structure can be recovered from the second moment, implying that only $n/\sigma^{4}\to\infty$ samples are necessary for accurate estimation in the high noise regime.
@@ 4.1 The second moment of cryo-EM
Using spherical coordinates $(r,\theta,\phi)$, it is typical to model a molecular structure $x\in L^{2}(\mathbb{R}^{3})$ by discretizing $x(r,\theta,\phi)$ with $R$ samples $r_{1},\ldots,r_{R}$, of the radial coordinates and bandlimiting the corresponding spherical functions $x(r_{i},\theta,\phi)$. This is a standard assumption in the cryo-EM literature, see for example, [8]. Mathematically, this means that we approximate the infinite-dimensional representation $L^{2}(\mathbb{R}^{3})$ with the finite-dimensional representation $V=\oplus_{\ell=0}^{L}V_{\ell}^{\oplus R}$, where $L$ is the bandlimit, and $V_{\ell}$ is the $(2\ell+1)$-dimensional irreducible representation of $\operatorname{SO}(3)$, corresponding to harmonic polynomials of frequency $\ell$ [^4] For convenience we consider a constant number of samples at the radial direction, namely, $R_{\ell}=R$ for all $\ell$, but omitting this assumption does not alter the analysis. In this model, an element $x\in V$ can be identified as an $R$-tuple $x=(x[1],\ldots,x[R])$, where
@@
where $X_{\ell}=\left(X_{\ell}^{m}[r_{i}]\right)_{m=-\ell,\ldots,\ell,i=1,\ldots R}$, contains the spherical harmonic coefficients of $x$. Therefore, the second moment determines the coefficient matrices $X_{\ell},\,\ell=0,\ldots,L$ up to the action of the ambiguity group $\prod_{\ell=0}^{L}U(2\ell+1)$. If we consider functions that are the Fourier transforms of real-valued functions on $\mathbb{R}^{3}$ (which is the case in cryo-EM), then the ambiguity group is $\prod_{\ell=0}^{L}O(2\ell+1)$.
@@ 4.2 Main result for cryo-EM
    If $K>M+2$, then a generic $x$ in a $\mathrm{O}(V)$-generic semi-algebraic set $\mathcal{M}$ of dimension $M$ is determined by its second moment, up to a sign. Likewise, $K>2M+2$ then every $x$ in a generic $\mathrm{O}(V)$-semi-lgebraic set $\mathcal{M}$ of dimension $M$ is determined by its second moment, up to a sign.
@@
In [15], the authors studied the cryo-EM model for signals having a sparse expansion with respect to a generic orthonormal basis. Specifically, the authors proved that if $\mathcal{V}$ is a generic orthonormal basis and the expansion of $x\in\mathbb{R}^{N}$ with respect to this basis is $M$-sparse (meaning at most $M$ basis coefficients are non-zero) with $\frac{M}{N}\approx\frac{1}{3}$, then a generic $x$ is determined (up to a sign) from its second moment. The set of $M$-sparse vectors with respect to a fixed orthonormal basis is a union of $\binom{N}{M}$ $M$-dimensional linear subspaces, and the set of $M$-sparse vectors with respect to a generic orthonormal basis $\mathcal{V}$ is an example of an $M$-dimensional $\mathrm{O}(V)$-generic semi-algebraic set. Notably, the bounds from Theorem 4.1 substantially beat the bound obtained in [15]. Indeed, for $\frac{M}{N}\approx\frac{2}{3}$ we can recover a generic $M$-sparse signal, and for $\frac{M}{N}\approx\frac{1}{3}$ our theorem implies that we can recover every $M$-sparse signal in a generic orthonormal basis.
@@
One of the prime motivations behind the cryo-EM technology is its potential to elucidate multiple conformations (shapes) of biological molecules [9]. Different conformations are associated with different functions of the molecule. Recovering the different conformations, referred to as the heterogeneity problem, is extremely important, giving rise to developing a wide variety of approaches and techniques. The heterogeneity problem is typically classified into two categories: discrete heterogeneity and continuous heterogeneity. In the former, the molecule may appear in several distinct and stable conformations. In the latter and more challenging setup, molecules may present a continuum of conformations, thereby exacerbating the computational challenge [46].
Our results state that if the 3-D structure lies in a low-dimensional semi-algebraic set, then it can be recovered from fewer observations (Corollary 4.2). This might be crucial for the discrete heterogeneity problem, where the total number of measurements is split among the different conformations. The situation is more complicated in the continuous heterogeneity problem. This problem is clearly ill-posed without prior on the distribution of conformations since each tomographic projection (measurement) might be associated with a different conformation. Our results suggest that semi-algebraic priors on the continuous space of conformations can render the continuous heterogeneity problem well-posed. Indeed, among the successful techniques proposed in recent years, two outstanding approaches are based on either assuming that the conformations lie on a linear subspace [4, 30] or can be represented using a deep generative model [53, 20]; both of these are semi-algebraic priors. It is important to note, however, that since the prior in the continuous heterogeneity case is not only on the 3-D structure but also on the distribution of conformations, the situation is more subtle than what we consider here.
@@ 4.3 Multi-reference alignment
where $g_{1},\ldots,g_{n}\in G$ are random elements drawn from a uniform distribution over the group $G$, and $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2}I)$ i.i.d. The goal is estimating $x$ from $y_{1},\ldots,y_{n}$, while the group elements are treated as nuisance variables. Evidently, the MRA observations are just noisy realizations of the signal recovery problem (3.1).
The multi-reference alignment model was first suggested as an abstraction of the cryo-EM model, where the tomographic projection is ignored [8]. Yet, in recent years this problem has been studied in more generality as a prototype of statistical models with intrinsic algebraic structures, e.g., [11, 7, 36]. As in the cryo-EM case, the sample complexity of the multi-reference alignment model is determined by the lowest order moment that identifies the signal [2], and it was shown that in many cases, the signal is determined only by the third moment [7]. Thus, $n/\sigma^{6}\to\infty$ is a necessary condition for recovery. Similar to the cryo-EM problem, this pessimistic result led to a series of papers, analyzing when a sparse signal can be recovered from the second moment (implying an improved sample complexity) [17, 29, 28, 15]. This paper generalizes these results as follows. Assume we acquire $n/\sigma^{4}\to\infty$ samples from (4.5). In this case, the empirical second moment of the observations approximates (almost surely) $\mathbb{E}[yy^{T}]$, up to a constant term that can be removed assuming the noise level $\sigma^{2}$ is known. Thus, a direct implication of Corollary 3.1 is the following result on the sample complexity of multi-reference alignment with a generic semi-algebraic prior.

### v3

@@ 3 Signal recovery from the second moment
Most applications we consider in this paper involve signal recovery from the second moment. Thus, we are interested in the transversality of semi-algebraic sets with orbits of the ambiguity group of the second moment. In this section, we first present the model of recovering a signal translated by random group actions from the second moment, and then specialize Theorem 2.2 and Theorem 2.3 to this case. Finally, we discuss two important applications: factoring Gram matrices and phase retrieval.
@@ 3.2 Main results for signal recovery from the second moment
We view $V$ as a representation of $H$, where the $\ell$-th component of $H$, $\mathrm{O}(N_{\ell})$, acts diagonally. This means that if we view an element of $V_{\ell}^{R_{\ell}}$ as a tuple $v_{\ell}=(v_{\ell}[1],v_{\ell}[2],\ldots,v_{\ell}[R_{\ell}])$ with $v_{\ell}[r]\in V_{\ell}$, then $g\cdot v_{\ell}=(g\cdot v_{\ell}[1],g\cdot v_{\ell}[2],\ldots g\cdot v_{\ell}[R_{\ell}])$. Then, the generic orbits of $H$ have dimension
@@
In Section 4 we discuss the implications of Corollary 3.1 to cryo-EM—the prime motivation of this paper. Before we do this, we briefly discuss two additional applications: factoring Gram matrices under semi-algebraic priors and phase retrieval.
@@ 3.3.1 Factoring Gram matrices under semi-algebraic priors
As a warm-up application, we consider the simple case above where $L=1$; i.e., $V=W^{\oplus R}$ with $W$ an $N$-dimensional irreducible representation of $H$. An element $X\in V$ is simply an $N\times R$ matrix, and the second moment is the Gram matrix $X^{T}X$. Generally, the Gram matrix determines $X$ up to multiplication by a matrix in $O(N)$. However, Corollary 3.1 implies that an $N\times R$ matrix can be determined from its Gram matrix if it lies in a suitably low-dimensional generic semi-algebraic subset of $\mathbb{R}^{N\times R}$.
@@
    If $X$ is a generic matrix in an $\mathrm{O}(V)$-generic semi-algebraic subset of dimension $M<NR-\left(\dim\mathrm{O}(N)-\dim\mathrm{O}(N-R)\right)-2$, then $X$ can be recovered (up to a sign for $\mathrm{GL}(V)$) from its Gram matrix. Likewise, if $2M<NR-\left(\dim\mathrm{O}(N)-\dim\mathrm{O}(N-R)\right)-2$, then every $X$ in an $\mathrm{O}(V)$-generic semi-algebraic subset can be recovered up to a sign from its Gram matrix.
@@ 3.3.2 Phase retrieval
The phase retrieval problem entails recovering a signal from its power spectrum: the magnitudes of its Fourier transform. Clearly, the power spectrum determines the signal up to a product of $N$ phases (unitary matrices of dimension 1) in Fourier space. This problem has numerous applications in optical imaging, X-ray crystallography and more, see [47, 11, 28, 32], and references therein. Semi-algebraic priors are essential in phase retrieval. For example, in X-ray crystallography, the goal is to recover the sparsely-spread atoms that form an image [27, 14, 15]. In addition, deep generative models have been proven effective in different branches of phase retrieval in recent years [49, 42, 37, 33, 22].
Explicitly, the goal in phase retrieval is to recover a signal $x\in\mathbb{R}^{N}$ from its power spectrum,
@@
where $\hat{x}[i]$ is the $i$-th entry of the discrete Fourier transform of the signal $x$. The analysis of [25, Section 6.2] shows that the power spectrum is equivalent to the second moment of $x$ with respect to the action of the dihedral group $D_{2N}$, which acts on $\mathbb{R}^{N}$ by circular translation and reflection. As a representation of $D_{2N}$, the vector space $\mathbb{R}^{N}$ decomposes as a sum of irreducibles as follows.
@@
    $\mathbb{R}^{N}=V_{0}\oplus V_{1}\oplus\ldots\oplus V_{N/2-1}\oplus V_{N/2},$
@@
    $\mathbb{R}^{N}=V_{0}\oplus V_{1}\oplus\ldots\oplus V_{(N-1)/2},$
@@
In particular, the dimension of the orbits of $H=\prod\mathrm{O}(V_{i})$ is N/21N/2-1 if $N$ is even and (N1)/2(N-1)/2 if $N$ is odd. Thus, we can invoke Corollary 3.1 with $K=N/2+1$ (NN even) and $K=(N+1)/2$ (NN odd) to obtain the following result.
@@
In [25], the authors considered phase retrieval for sparse vectors in generic bases. There, by taking advantage of the fact that the set of sparse vectors is the union of linear subspaces and using algebro-geometric arguments, the bounds $N\geq 2M-1$ (generic signal recovery) and $N\geq 4M-3$ (all signal recovery) are obtained.
@@ 4 Cryo-EM
where $T$ is a tomographic projection $Tx(z_{1},z_{2})=\int_{\mathbb{R}}x(z_{1},z_{2},z_{3})dz_{3}$, $\varepsilon_{i}$ is a “noise” term drawn from an i.i.d. normal distribution with zero mean and variance $\sigma^{2}$, and $R_{\omega}$ rotates the function $x$, where $\omega$ is drawn from uniform distribution over $\operatorname{SO}(3).$ Importantly, the goal of the cryo-EM inverse problem is only to estimate the unknown function $x$ from $y_{1},\ldots,y_{n}$, while the unknown 3-D rotations $\omega_{1},\ldots,\omega_{n}$ are treated as nuisance variables.
One of the main challenges in processing cryo-EM data sets and analyzing their performance is the extremely high noise level. This is due to the limited number of electrons the microscope can transmit without damaging the biological sample. From an information-theoretic perspective, it was shown that in the high-noise regime $\sigma\to\infty$, the minimal number of samples required for accurate recovery, regardless of any specific algorithm, needs to scale as $n/\sigma^{2d}\to\infty$, where $d$ is the lowest order moment of the observations that determine the sought structure uniquely [2] [^4] This result is true when the dimension of the signal is finite, as we assume in this paper. In the very high-dimensional regime, the sample complexity is governed by other factors [44].. In [7], the authors gave computational evidence that the third-order moment determines generic signals while the second moment does not, implying that the minimal number of observations $n$ for generic signals scales rapidly with the noise level as $\sigma^{6}$. This motivates identifying classes of signals that can be determined from the second moment, and thus with fewer samples. For example, recent works showed that if the structure can be represented with only a few coefficients under some basis [16], or as a sparse mixture of Gaussians [17], then the signal can be recovered from the second moment. In this work, we consider the more general family of signals that lie in a semi-algebraic set (including sparse signals as in [16]) and derive conditions under which a molecular structure can be recovered from the second moment, implying that only $n/\sigma^{4}\to\infty$ samples are necessary for accurate estimation in the high noise regime.
@@ 4.1 The second moment of cryo-EM
Using spherical coordinates $(r,\theta,\phi)$, it is typical to model a molecular structure $x\in L^{2}(\mathbb{R}^{3})$ by discretizing $x(r,\theta,\phi)$ with $R$ samples $r_{1},\ldots,r_{R}$, of the radial coordinates and bandlimiting the corresponding spherical functions $x(r_{i},\theta,\phi)$. This is a standard assumption in the cryo-EM literature, see for example, [8]. Mathematically, this means that we approximate the infinite-dimensional representation $L^{2}(\mathbb{R}^{3})$ with the finite-dimensional representation $V=\oplus_{\ell=0}^{L}V_{\ell}^{\oplus R}$, where $L$ is the bandlimit, and $V_{\ell}$ is the $(2\ell+1)$-dimensional irreducible representation of $\operatorname{SO}(3)$, corresponding to harmonic polynomials of frequency $\ell$. [^5] For convenience, we consider a constant number of samples at the radial direction, namely, $R_{\ell}=R$ for all $\ell$, but omitting this assumption does not alter the analysis. In this model, an element $x\in V$ can be identified as an $R$-tuple $x=(x[1],\ldots,x[R])$, where
@@
where $X_{\ell}=\left(X_{\ell}^{m}[r_{i}]\right)_{m=-\ell,\ldots,\ell,i=1,\ldots R}$, contains the spherical harmonic coefficients of xx666 We note that when the distribution over $\operatorname{SO}(3)$ is non-uniform, the second moment takes a different form that depends explicitly on the distribution; see [46]. Specifically, in this case, the second moment is given by $\int_{\omega}\rho(\omega)(T(R_{\omega}\cdot x))\,(T(R_{\omega}\cdot x))^{*}\,d\omega,$ which defines a rich family of invariant functions of total degree three on $R(G)\times V\times V$, where $R(G)$ denotes the regular representation of $G$. The analysis presented in this paper does not address this more general setting.. Therefore, the second moment determines the coefficient matrices $X_{\ell},\,\ell=0,\ldots,L$ up to the action of the ambiguity group $\prod_{\ell=0}^{L}U(2\ell+1)$. If we consider functions that are the Fourier transforms of real-valued functions on $\mathbb{R}^{3}$ (which is the case in cryo-EM), then the ambiguity group is $\prod_{\ell=0}^{L}O(2\ell+1)$.
@@ 4.2 Main result for cryo-EM
    If $K>M+2$, then a generic $x$ in a $\mathrm{O}(V)$-generic semi-algebraic set $\mathcal{M}$ of dimension $M$ is determined by its second moment, up to a sign. Likewise, $K>2M+2$ then every $x$ in a generic $\mathrm{O}(V)$-semi- algebraic set $\mathcal{M}$ of dimension $M$ is determined by its second moment, up to a sign.
@@
We note that the sample complexity results in the MRA literature are asymptotic in nature, focusing primarily on the rate at which the number of observations must grow (in this case, faster than $\sigma^{4}$). Nevertheless, there is substantial numerical evidence, dating back more than 25 years, that these asymptotic predictions accurately reflect the model’s behavior in the finite-sample regime; see, for example, [48, 12, 2].
@@
In [16], the authors studied the cryo-EM model for signals having a sparse expansion with respect to a generic orthonormal basis. Specifically, the authors proved that if $\mathcal{V}$ is a generic orthonormal basis, and the expansion of $x\in\mathbb{R}^{N}$ with respect to this basis is $M$-sparse (meaning at most $M$ basis coefficients are non-zero) with $\frac{M}{N}\approx\frac{1}{3}$, then a generic $x$ is determined (up to a sign) from its second moment. The set of $M$-sparse vectors with respect to a fixed orthonormal basis is a union of $\binom{N}{M}$ $M$-dimensional linear subspaces, and the set of $M$-sparse vectors with respect to a generic orthonormal basis $\mathcal{V}$ is an example of an $M$-dimensional $\mathrm{O}(V)$-generic semi-algebraic set. Notably, the bounds from Theorem 4.1 substantially beat the bound obtained in [16]. Indeed, for $\frac{M}{N}\approx\frac{2}{3}$ we can recover a generic $M$-sparse signal, and for $\frac{M}{N}\approx\frac{1}{3}$ our theorem implies that we can recover every $M$-sparse signal in a generic orthonormal basis.
@@
One of the prime motivations behind the cryo-EM technology is its potential to elucidate multiple conformations (shapes) of biological molecules [10]. Different conformations are associated with different functions of the molecule. Recovering the different conformations, referred to as the heterogeneity problem, is extremely important, giving rise to the development of a wide variety of approaches and techniques. The heterogeneity problem is typically classified into two categories: discrete heterogeneity and continuous heterogeneity. In the former, the molecule may appear in several distinct and stable conformations. In the latter and more challenging setup, molecules may present a continuum of conformations, thereby exacerbating the computational challenge [51].
Our results state that if the 3-D structure lies in a low-dimensional semi-algebraic set, then it can be recovered from fewer observations (Corollary 4.2). This might be crucial for the discrete heterogeneity problem, where the total number of measurements is split among the different conformations. The situation becomes more complicated in the context of continuous heterogeneity. This problem is clearly ill-posed without prior on the distribution of conformations, since each tomographic projection (measurement) might be associated with a different conformation. Our results suggest that semi-algebraic priors on the continuous space of conformations can render the continuous heterogeneity problem well-posed. Indeed, among the successful techniques proposed in recent years, two outstanding approaches are based on either assuming that the conformations lie on a linear subspace [4, 31] or can be represented using a deep generative model [58, 21]; both of these are semi-algebraic priors. It is important to note, however, that since the prior in the continuous heterogeneity case is not only on the 3-D structure but also on the distribution of conformations, the situation is more subtle than what we consider here.
@@ 4.3 Multi-reference alignment
where $g_{1},\ldots,g_{n}\in G$ are random elements drawn from a uniform distribution over the group $G$, and $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2}I)$ i.i.d. is the noise. The goal is to estimate $x$ from $y_{1},\ldots,y_{n}$, while the group elements are treated as nuisance variables. Evidently, the observations are just noisy realizations of the signal recovery problem (3.1).
The multi-reference alignment model was first suggested as an abstraction of the cryo-EM model, where the tomographic projection is ignored [8]. Yet, in recent years, this problem has been studied in more generality as a prototype of statistical models with intrinsic algebraic structures, e.g., [12, 7, 38]. As in the cryo-EM case, the sample complexity of the multi-reference alignment model is determined by the lowest order moment that identifies the signal [2], and it was shown that in many cases, the signal is determined only by the third moment [7]. Thus, $n/\sigma^{6}\to\infty$ is a necessary condition for recovery. Similar to the cryo-EM problem, this pessimistic result led to a series of papers analyzing when a sparse signal can be recovered from the second moment (implying an improved sample complexity) [18, 30, 29, 16]. This paper generalizes these results as follows. Assume we acquire $n/\sigma^{4}\to\infty$ samples from (4.5). In this case, the empirical second moment of the observations approximates (almost surely) $\mathbb{E}[yy^{T}]$, up to a constant term that can be removed assuming the noise level $\sigma^{2}$ is known. Thus, a direct implication of Corollary 3.1 is the following result on the sample complexity of multi-reference alignment with a generic semi-algebraic prior.